OpenStudy (anonymous):

can i get help solving f(x)=10-x, find inverse f-1(x)_______

OpenStudy (anonymous):

First of all you must know that not all functions have inverses. Therefore to verify you must us the following statement: if \[f(a)=f(b)\]then \[a=b\] Now just plug in the statement and you have that \[10 - a = 10- b\] which by simple arithmetic you have that \[a = b\] which states that the function has an inverse. Now to find the inverse you just substitute \[x\] by \[f^{-1}(x)\] and \[f(x)\] by \[x\] which in this case is the following:\[x = 10 - f ^{-1}(x)\] and clear for the inverse function:\[x - 10 = -f ^{-1}(x)\] ; \[-x +10 = f ^{-1}(x)\] which is the result you needed. Hope it helps form the beautiful island of Puerto Rico. Peace, Love and Happiness

OpenStudy (anonymous):

thanks